# EmoGist: Efficient In-Context Learning for Visual Emotion Understanding

### v1

@@ 5 Conclusion and Future Work
In this paper, we introduced EmoGist, a training-free in-cotext learning method for visual emotion understanding with large vision-language models. We observe a significant amount of improvements over the zero-shot and naive ICL baselines across SOTA LVLMs of 2 and 7 billion parameters. In particular, we find that $\text{EmoGist}_{\text{e}}$, the variant of our method with a simple ensembling, achieves robust performance improvements with higher precision. In future work, we’d like to explore more deeply into the label explanations generated by the strong LVLMs and investigate various reasoning strategies for obtaining emotion label descriptions that are more transferrable across different domains.
@@
Due to limited computational resources available to the authors, we perform our experiments on a limited subset of publicly available large vision-language models with 2 and 7 billion parameters. While we anticipate that our findings would hold overall for other model sizes, we do not provide any direct evidence. We also note that we only tried one model each for the embedding model and the explanation generation LVLM, as stated in section 3.
@@ Appendix A Dataset Information
For each dataset, we treat the training set as the pool of example images for explanation generation, use the validation set for hyperparameter tuning, and test our method and baselines against the test set.
@@ A.1 Memotion
Appendix B Hardwares and Softwares Used
To fully utilize all the GPU resources available to us, we ran all of test predictions using multiple NVIDIA V100, GTX TITAN Xp and GTX TITAN X GPUs.
@@ Appendix B Hardwares and Softwares Used
For our label explanation generation, we used a Mac Studio hardware with M1 Ultra CPU and 128GB of RAM, using the version 0.25.0 of mlx (Hannun et al., 2023) and 0.1.23 of mlx-vlm.
Appendix C Prompts used for label description generation
@@
These are the examples of images with the sentiment of "{s_label}". Based on these examples, what are the common features of images to be felt "{s_label}" by its viewers? Do not reference example images directly.
Appendix D Prompts used for classification
D.1 Multi-label
[…]
D.2 Multi-class
[…]

### v2

@@ 5 Conclusion and Future Work
In this paper, we introduced EmoGist, a training-free in-cotext learning method for visual emotion understanding with large vision-language models. We observe a significant amount of improvements over the zero-shot and naive ICL baselines across SOTA LVLMs of 2 and 7 billion parameters. In particular, we find that $\text{EmoGist}_{\text{e}}$, the variant of our method with simple ensembling, achieves robust performance improvements with higher precision. In future work, we’d like to explore more deeply into the label descriptions generated by the strong LVLMs and investigate various reasoning strategies for obtaining emotion label descriptions that are more transferrable across different domains.
Acknowledgments
The work was supported by NSF CAREER award IIS2048001 and the DARPA CCU program. Contents do not necessarily represent the official views of, nor an endorsement by, DARPA, or the US Government.
@@ Acknowledgments
Due to limited computational resources available to the authors, we perform our experiments on a limited subset of publicly available large vision-language models with 2 and 7 billion parameters. While we anticipate that our findings would hold overall for other model sizes, we do not provide any direct evidence. We also note that we only tried one model each for the embedding model and the description generation LVLM, as stated in section 3.
@@ Appendix A Dataset Information
For each dataset, we treat the training set as the pool of example images for description generation, use the validation set for hyperparameter tuning, and test our method and baselines against the test set.
@@ A.1 Memotion
Appendix B Experimental Settings
B.1 Hardwares and Softwares Used
To fully utilize all the GPU resources available to us, we ran all of our validation and test predictions, and embedding generations using multiple NVIDIA V100, GTX TITAN Xp and GTX TITAN X GPUs.
@@ B.1 Hardwares and Softwares Used
For our label description generation, we used a Mac Studio hardware with M1 Ultra CPU and 128GB of RAM, using the version 0.25.0 of mlx (Hannun et al., 2023) and 0.1.23 of mlx-vlm.
B.2 Random Seeds
For the results shown in Table 1 and Table 2, we run each test for 6 different random seeds: 21, 42, 63, 84, 105, 126.
Appendix C Sensitivity analysis and hyperparameter tuning for the number of clusters $k$
We determine the optimal number of clusters for each dataset and test-time LVLM combination. Initially, we apply $k$-means clustering to both FI and Memotion datasets for $k$ values of 2, 4, and 6. Subsequently, we generate cluster-based label descriptions for each $k$ using the Qwen2.5-VL 72B model. We then classify the validation set based on these clusters and text descriptions, 3 times for each 7B/2B LVLM using the following random seeds: 21, 42, 63. The optimal $k$ is ultimately identified by identifying $k$ that achieved the most number of highest validation F1 scores across the three seeds. If there’s no winner, we ran the validation for additional seeds (84, 105, 126) until the clear winner is found.
Please see Figure 9 and Figure 10 for the validation set results of $\text{EmoGist}_{\text{n}}$ and $\text{EmoGist}_{\text{e}}$ with 7B models on FI, Figure 11 and Figure 12 for the validation set results of $\text{EmoGist}_{\text{n}}$ and $\text{EmoGist}_{\text{e}}$ with 7B models on Memotion, and Figure 13 and Figure 14 for the validation set results of $\text{EmoGist}_{\text{e}}$ with 2B models on FI and Memotion.
Appendix D Prompts used for label description generation
**Images from the cluster go here**
@@ Appendix D Prompts used for label description generation
These are the examples of images with the sentiment of “{s_label}". Based on these examples, what are the common features of images to be felt “{s_label}" by its viewers? Do not reference example images directly.
Figure 3: Prompts used for label description generation.
Appendix E Prompts used for zero-shot and EmoGist
**Test image goes here**
**For EmoGist, label description goes here**
Question: Does this image match the sentiment label ‘{s_label}’? Answer:
Answer with ‘Yes’ or ‘No’.
Figure 4: Prompts used for multi-label zero-shot classification and EmoGist.
**Test image goes here**
**For EmoGist, label description goes here**
Question: Which of the sentiment labels in the following list does this image belong to? List: [“amusement", “anger", “awe", “contentment", “disgust", “excitement", “fear", “sadness"] Answer:
Answer with the exact sentiment label as it appears in the list.
Figure 5: Prompts used for multi-class zero-shot classification and EmoGist.
Appendix F Prompts used for ICL baselines
**ICL example images 1 to 4 goes here**
**Test image goes here**
## Image 1
Question: Does this image matches the sentiment label ’{s_label}’? Answer: Yes
## Image 2
Question: Does this image matches the sentiment label ’{s_label}’? Answer: Yes
## Image 3
Question: Does this image matches the sentiment label ’{s_label}’? Answer: Yes
## Image 4
Question: Does this image matches the sentiment label ’{s_label}’? Answer: Yes
## Image 5
Question: Does this image matches the sentiment label ’{s_label}’? Answer:
Answer with ‘Yes’ or ‘No’.
Figure 6: Prompts used for multi-label $\text{ICL}_{\text{sim}}$.
**ICL example images 1 to 4 goes here**
**Test image goes here**
## Image 1
Question: Which of the sentiment labels in the following list does this image belong to? List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: {s_label}
## Image 2
Question: Which of the sentiment labels in the following list does this image belong to? List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: {s_label}
## Image 3
Question: Which of the sentiment labels in the following list does this image belong to? List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: {s_label}
## Image 4
Question: Which of the sentiment labels in the following list does this image belong to? List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: {s_label}
## Image 5
Question: Which of the sentiment labels in the following list does this image belong to? List: [“amusement", “anger", “awe", “contentment", “disgust", “excitement", “fear", “sadness"] Answer:
Answer with the exact sentiment label as it appears in the list.
Figure 7: Prompts used for multi-class $\text{ICL}_{\text{sim}}$.
**ICL example images 1 to 8 goes here**
**Test image goes here**
## Image 1
Question: Which of the sentiment labels in the following list does this image belong to? List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: amusement
## Image 2
Question: Which of the sentiment labels in the following list does this image belong to? List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: anger
## Image 3
Question: Which of the sentiment labels in the following list does this image belong to? List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: awe
## Image 4
Question: Which of the sentiment labels in the following list does this image belong to? List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: contentment
## Image 5
Question: Which of the sentiment labels in the following list does this image belong to? List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: disgust
## Image 6
Question: Which of the sentiment labels in the following list does this image belong to? List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: excitement
## Image 7
Question: Which of the sentiment labels in the following list does this image belong to? List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: fear
## Image 8
Question: Which of the sentiment labels in the following list does this image belong to? List: [‘‘amusement", ‘‘anger", ‘‘awe", ‘‘contentment", ‘‘disgust", ‘‘excitement", ‘‘fear", ‘‘sadness"] Answer: sadness
## Image 9
Question: Which of the sentiment labels in the following list does this image belong to? List: [“amusement", “anger", “awe", “contentment", “disgust", “excitement", “fear", “sadness"] Answer:
Answer with the exact sentiment label as it appears in the list.
Figure 8: Prompts used for multi-class, $\text{ICL}_{\text{all}}$.
Figure 9: Validation set results for $\text{EmoGist}_{\text{n}}$ with 7B models on FI.
Figure 10: Validation set results for $\text{EmoGist}_{\text{e}}$ with 7B models on FI.
Figure 11: Validation set results for $\text{EmoGist}_{\text{n}}$ with 7B models on Memotion.
Figure 12: Validation set results for $\text{EmoGist}_{\text{e}}$ with 7B models on Memotion.
Figure 13: Validation set results for $\text{EmoGist}_{\text{e}}$ with 2B models on FI.
Figure 14: Validation set results for $\text{EmoGist}_{\text{e}}$ with 2B models on Memotion.